OpenStudy (anonymous):

How do you find the area of an acute triangle when you're only given the side lengths????

OpenStudy (amistre64):

one way is; you can use the law of cosines to find an angle and use that angle in the side*side*sin(angle)/2 formula

OpenStudy (amistre64):

the other is to use herons formula

OpenStudy (anonymous):

so the area = \[a timesb \times \sin \] of the angle contained between a and b?

OpenStudy (anonymous):

then, divided by two

OpenStudy (amistre64):

thats right; so if you can find the angle between two sides, you can find the area

OpenStudy (anonymous):

okay! Thanks :)

OpenStudy (amistre64):

in my picture; base = b; height = a sin(angle) which fits into the normal Area = b*h/2 that you are familiar with